# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v2

@@ Abstract
Machine Learning, ICML
{fengc, aravento, sganguli, shauld}@stanford.edu, nancheng@umich.edu
@@ 1 Introduction
Scaling test-time compute has been integral to unprecedented improvements in LLMs’ reasoning skills for complex tasks such as math and coding. Thus, test-time compute has emerged as a new dimension for improving LLMs, leading to a key tradeoff between allocating additional compute to inference versus pretraining (Snell et al., 2024). Diverse test-time strategies include Chain-of-Thought (CoT) (Wei et al., 2022), tree-of-thought (Yao et al., 2023), self-consistency (Wang et al., 2023), self-reflection (Shinn et al., 2023), self-critique (Saunders et al., 2022), self-verification (Weng et al., 2023) and Monte-Carlo tree search (Zhao et al., 2023). These have shown great success in boosting model performance in the post-training phase or at inference time. More recently, OpenAI’s O1 model (OpenAI, 2024) and DeepSeek’s R1 model (DeepSeek-AI: Daya Guo et al., 2025) have combined some of these strategies with reinforcement learning to generate high-quality reasoning traces for problems of various difficulty levels, demonstrating clear performance improvements as more test-time compute is allocated.
These successes fit into a broader paradigm in which a frontier model is first fine-tuned on a reasoning task with supervised fine-tuning (SFT) (Wei et al., 2022; Ouyang et al., 2022; Chung et al., 2022), and then a test-time algorithm is applied to model outputs or reasoning traces to improve performance (Yao et al., 2023; Wang et al., 2023; Chen et al., 2021). Many test-time algorithms are independent of the fine-tuning process. As a result, the fine-tuning is agnostic to and thus decoupled from the test-time algorithm (Chow et al., 2024). However, for a given choice of test-time strategy and compute budget, it is not a priori clear which fine-tuning approach, including the loss objective, would be best aligned with the test-time strategy so as to maximize the test accuracy under the overall strategy.
Our work studies the problem of aligning fine-tuning and test-time algorithms. We consider what is perhaps the simplest setting, supervised fine-tuning with CE loss under the pass@N test-time strategy. This setting reveals a case of misalignment: standard SFT is not the right choice for maximizing performance under pass@N. We believe that this kind of misalignment presents itself in several combinations of fine-tuning/test-time approaches, motivating our thorough study in this paper. Our main contributions are,
@@
  • We propose an alternate loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers in MATH and MiniF2F (Sections 5.1, 5.2 and 5.3).
@@ 2 Related Works
Test-time compute and pass@N strategy. A growing body of work explores how training and test-time strategies interact to affect model performance. Jones (2021) highlights a tradeoff between train and test-time compute in a controlled board game setting. Closely related to our work, Brown et al. (2024) identify an exponentiated power law between coverage and the number of in-context samples. Snell et al. (2024), in turn, explore compute-optimal strategies for scaling test-time compute. Our paper focuses primarily on the pass@N test-time strategy. Gui et al. (2024) show that, in the best-of-N setting, the sample distribution is nearly optimally aligned with the reward model. To mitigate the excessive test-time compute costs associated with best-of-N sampling, Gui et al. (2024); Sessa et al. (2024) propose alignment algorithms to distill the sample distribution. Li et al. (2024) argue that fine-tuning with cross-entropy loss can limit output diversity and propose a maximum entropy-based method to address this issue. Our work similarly identifies limitations of cross-entropy when scaling test-time compute, but focuses on the alignment of training and test-time strategies. Zhang et al. (2024) explore methods for preventing catastrophic forgetting in supervised fine-tuning over diverse domains.
@@
Post-training for mathematical reasoning. Various post-training techniques have been proposed to improve mathematical reasoning in LLMs. Instruction-tuning and reinforcement learning with human feedback have been shown to boost model performance on math (Yue et al., 2024; Lightman et al., 2024; Uesato et al., 2022), while continued training on math- or code-specific domain data enhances models’ reasoning abilities for downstream mathematical tasks (Lewkowycz et al., 2022; Azerbayev et al., 2024; Yang et al., 2024; Shao et al., 2024; Ying et al., 2024). Rejection-sampling (Zelikman et al., 2022) and self-improvement techniques (Qi et al., 2024), in turn, are useful to augment the training data for SFT. More recent approaches (OpenAI, 2024; DeepSeek-AI: Daya Guo et al., 2025) have incorporated reinforcement learning and achieved exceptional reasoning capabilities in various domains, such as math and coding. Although our paper primarily focuses on supervised fine-tuning to enhance pretrained models’ math capabilities, our loss function can be applied to other settings that train under CE loss, such as continual training, instruction-tuning, and data augmentation.
@@
Data pruning and hard example mining. The interpretation of the loss function we derive below can be related to data pruning and hard example mining (see discussion of sample difficulty in Section 4.4). Data selection is often applied to curate high-quality datasets for pretraining (Marion et al., 2023), where Sorscher et al. (2022) shows that pruning easy samples can improve pretraining loss scaling as a function of dataset size. Zhou et al. (2023); Ye et al. (2025) demonstrate that supervised fine-tuning on even a very small dataset can yield remarkably strong performance in alignment and reasoning tasks. On the other hand, hard example mining focuses on identifying and emphasizing challenging samples to improve model performance (Shrivastava et al., 2016). In the domain of mathematical reasoning, Tong et al. (2024) find a difficulty imbalance in rejection-sampled datasets and show that more extensive training on difficult samples improves model performance.
Our paper is closely related to a concurrent paper by Chow et al. (2024), which derives a similar training objective for RL to directly optimize for the best-of-N test-time strategy.
@@ 3 Problem setup
Given a vocabulary set $W$, we consider a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{M}$, where $x^{(i)}\in W^{n_{i}}$ is a prompt, $y^{(i)}\in W^{m_{i}}$ is its ground-truth completion, and $n_{i}$ and $m_{i}$ are the prompt and completion lengths. In the context of math, $x^{(i)}$ is the problem statement and $y^{(i)}$ is its solution. To model the conditional distribution $p(y^{(i)}|x^{(i)})$ we use an autoregressive transformer model (Vaswani et al., 2017), which is traditionally trained by minimizing the cross-entropy loss
@@
To use and evaluate the model at test time, we assume the existence of an efficient oracle verifier $V$ which takes as input an $(x,y)$ pair and returns $V(x,y)=1$ if $y$ is a correct completion of $x$ and otherwise returns $V(x,y)=0$. Practical examples of verifiers include compilers or pre-defined unit tests for coding problems, or automatic proof checkers in mathematical theorem proving. In such applications, a simple method, known as pass@N, for trading test-time compute for accuracy involves sampling $N$ completions from $\hat{p}$ given the test prompt $x$ and applying the verifier $V$ to all of them to search for a correct solution. The probability of a correct answer is then no longer the probability that $1$ completion is correct, but rather the probability that at least one of $N$ is correct. This probability, for a dataset $\mathcal{D}$, is given by the pass@N coverage metric
$\mathcal{C}_{\mathcal{D}}^{N}=\mathop{\mathbb{E}}_{\begin{subarray}{c}x\sim\mathcal{D}\\ \{y_{i}\}_{i\in[N]}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\hat{p}(\cdot|x)\end{subarray}}\mathbb{P}(\exists j\in[N]\,s.\,t.\,V(x,y_{j})=1).$ (2)
@@ 4.1 The CE loss induces overfitting for pass@N
Table 1: Pass@N coverage metric on the MATH test set for a Llama-3-8B-base model fine-tuned with CE loss on direct answers from the MATH training set. Surprisingly, Pass@N test accuracy at large $N$ decreases with number of training epochs.
@@
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than LLama-3-8B-Instruct to avoid potential leakage of the MATH dataset into LLama-3-8B-Instruct through post-training. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer, without a reasoning trace. We will discuss training with CoT in Section 5.4 for MATH reasoning traces. We confirm that the results presented in this section also hold for a larger model, Llama-3-70B-base, and refer readers to Section C.2 for additional results.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
However, Lemma 4.1 can fail when there is more than one problem. To understand this in a simple setting, consider a test set with two problems $x_{1}$ and $x_{2}$ with corresponding unique correct answers $y_{1}$ and $y_{2}$. Consider two models. The first model assigns probabilities $\hat{p}_{1}(y_{1}|x_{1})=1$ and $\hat{p}_{1}(y_{2}|x_{2})=0$. If we think of the probability a model assigns to an answer as its confidence in that answer, this model is highly confident and correct for problem $x_{1}$, but highly confident and wrong for $x_{2}$. Its pass@1 coverage is thus 50%. Moreover, since it always gets $x_{1}$ right and $x_{2}$ wrong, its pass@N coverage remains at 50%. In contrast, consider a second model which assigns probabilities $\hat{p}_{2}(y_{1}|x_{1})=\hat{p}_{2}(y_{2}|x_{2})=0.1$. This model has high confidence (0.9) but unfortunately on incorrect answers. Therefore its pass@1 accuracy is only 10%. However, it is willing to explore or hedge by placing some low confidence ($0.1$) on the other answer, which happens to be correct. Thus if this model samples $N$ times, the pass@N coverage increases with $N$, eventually approaching 100% as $N\rightarrow\infty$. Thus the first model outperforms the second in terms of pass@1 but not pass@N for large $N$. This indicates a tradeoff amongst policies: those that do better on pass@1 may not do better at pass@N. Of course the best one can do is be confident and correct, corresponding to the optimal model with $\hat{p}^{*}(y_{1}|x_{1})=\hat{p}^{*}(y_{2}|x_{2})=1$. However, this toy example reveals that if one cannot guarantee correctness, it can be beneficial to limit confidence and explore more solutions, especially if one can sample at large $N$.
@@
This lower bound is a monotonically decreasing function of $N$ (Figure 13), implying that at small $N$, optimal policies for pass@N must have high max confidence, thereby favoring exploitation and discouraging exploration. Note in the limit $N\to 1^{+}$, the lower bound is always $1$, recovering the intuitive result that the optimal policy for pass@1 is to place 100% confidence on the highest accuracy answer.
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
Figure 1: A model trained with CE loss becomes overconfident in its greedy completions, which harms its pass@N coverage; our proposed DCO objective limits this overconfidence. We fine-tune a Llama-3-8B base model on the MATH dataset to produce direct answers without a reasoning trace. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when sampling autoregressively and choosing the most likely token at each step. Leftmost: The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. Left: At the end of the training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up $N$. Right: Same as leftmost but shown for the DCO loss with $N=256$. Relative to the CE loss, the model trained on DCO shows a much milder overconfidence effect. Rightmost: The confidence distribution of the greedy completions after four epochs with DCO for various choices of $N$. As $N$ increases, the model’s confidence on the greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
@@
To summarize the consequences of our lemmas above, among the space of approximately well calibrated policies in which higher model confidence on an answer is correlated with higher accuracy on the answer, Lemma 4.2 suggests that at large $N$ it is beneficial to unconfidently explore by assigning low model confidence to many answers, while Lemma 4.3 suggests that at small $N$ it is beneficial to confidently exploit by assigning high model confidence to one or a few answers. These lemmas make a prediction that could explain the empirical observation in Table 1 that pass@N test performance for large $N$ degrades over epochs when pass@1 is maximized at training time: namely, maximization of pass@1 makes the model overconfident, thereby preventing good performance on pass@N.
@@
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas 4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure 1 (leftmost). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure 1 (left) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance. In Section C.3, we observe a similar trend of overconfidence on the out-of-distribution test set AIME24, further highlighting the generality of this failure mode.
4.4 Easy data drives overconfidence
We have shown above that overconfidence limits performance gains from scaling test-time compute. Here we investigate whether all training data contribute equally to the drop in pass@N test performance. To quantify this effect, we define the test loss as the negative log-probability of coverage, $\mathcal{L}_{\text{test}}^{N}=-\sum_{(x,y)\in\mathcal{D}^{\mathrm{test}}}\mathcal{\log}\ \mathcal{C}_{(x,y)}^{N}$, where $\mathcal{C}_{(x,y)}^{N}$ denotes the coverage for a single test example $(x,y)$. To measure how a training batch $\mathcal{B}$ influences the final test loss, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ for the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$, where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example. A negative value, $\nabla_{g}\mathcal{L}_{\text{test}}^{N}<0$, indicates that a gradient step in the direction of $g$ decreases the test loss.

### v3

@@ 1 Introduction
Scaling test-time compute has been integral to unprecedented improvements in LLMs’ reasoning skills for complex tasks such as math and coding. Thus, test-time compute has emerged as a new dimension for improving LLMs, leading to a key tradeoff between allocating additional compute to inference versus pretraining [1]. Diverse test-time strategies include Chain-of-Thought (CoT) [2], tree-of-thought [3], self-consistency [4], self-reflection [5], self-critique [6], self-verification [7] and Monte-Carlo tree search [8]. These have shown great success in boosting model performance in the post-training phase or at inference time. More recently, OpenAI’s O1 model [9] and DeepSeek’s R1 model [10] have combined some of these strategies with reinforcement learning to generate high-quality reasoning traces for problems of various difficulty levels, demonstrating clear performance improvements as more test-time compute is allocated.
These successes fit into a broader paradigm in which a frontier model is first fine-tuned on a reasoning task with supervised fine-tuning (SFT) [2, 11, 12], and then a test-time algorithm is applied to model outputs or reasoning traces to improve performance [3, 4, 13]. Many test-time algorithms are independent of the fine-tuning process. As a result, the fine-tuning is agnostic to and thus decoupled from the test-time algorithm [14]. However, for a given choice of test-time strategy and compute budget, it is not a priori clear which fine-tuning approach, including the loss objective, would be best aligned with the test-time strategy so as to maximize the test accuracy under the overall strategy.
Our work studies the problem of aligning fine-tuning with test-time algorithms. We consider what is perhaps the simplest setting, SFT with CE loss and pass@N as the test-time strategy. This setting reveals a case of misalignment: standard SFT is not maximizing performance under pass@N. We believe that this misalignment presents itself in several combinations of fine-tuning/test-time approaches, motivating our thorough study in this paper. Our main contributions are,
@@
  • We propose a loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers on MATH and MiniF2F (Sections 5.1, 5.2 and 5.3).
@@ 2 Related Works
Test-time compute and pass@N strategy. A growing body of work explores how training and test-time strategies interact to shape model performance. Jones [15] highlight a tradeoff between train and test-time compute in a controlled board game setting. Brown et al. [16] identify an exponentiated power law between coverage and number of in-context samples, while Snell et al. [1] explore compute-optimal strategies for scaling test-time inference. Our paper focuses primarily on the pass@N test-time strategy. Gui et al. [17] show that, in the best-of-N setting, the sample distribution is well-aligned with the reward model, and alignment methods have been proposed to distill this distribution and reduce sampling costs [17, 18]. Closely related to our work, Chow et al. [14] derives a similar training objective for reinforcement learning (RL) to directly optimize for the best-of-N test-time strategy. Li et al. [19] argue that fine-tuning with cross-entropy loss limits output diversity and propose a maximum entropy-based method to address this issue. Our work similarly identifies limitations of cross-entropy when scaling test-time compute, but focuses on the alignment of training and test-time strategies. In addition, Yue et al. [20] argue that fine-tuning with Group Relative Policy Optimization (GRPO) does not improve reasoning capabilities beyond the base model when evaluated with pass@N strategy. We will show that GRPO training may also lead to overconfidence in Figure 3.
@@
Post-training for mathematical reasoning. Various post-training techniques have been developed to improve mathematical reasoning in LLMs. Instruction-tuning and reinforcement learning with human feedback boost model performance on math tasks [21, 22, 23], while continued training on domain-specific math or code data also boosts downstream reasoning [24, 25, 26, 27, 28]. Rejection-sampling [29] and self-improvement [30] methods augment training data for SFT. Recent approaches [9, 10] have incorporated RL to achieve exceptional reasoning capabilities. GRPO [27, 10] is a newly proposed RL algorithm demonstrating strong improvements in mathematical reasoning. Several recent studies have further addressed response length and problem difficulty biases [31], as well as conducted extensive analyses of accuracy and response length across a wide variety of base models [31, 32]. While our primary focus is SFT, our loss function can be applied to other settings that train under CE loss, such as continual training, instruction-tuning, and data augmentation.
@@
Data pruning and hard example mining. The loss function we derive below can be viewed from the lens of data pruning and hard example mining. Data selection is used to curate high-quality datasets for pretraining [33], where Sorscher et al. [34] show that pruning easy samples can improve loss scaling as a function of dataset size. Zhou et al. [35], Ye et al. [36] demonstrate that, in SFT, even small datasets can yield strong alignment and reasoning performance. Hard example mining focuses on identifying and emphasizing challenging samples to improve model performance [37]. In the domain of mathematical reasoning, Tong et al. [38] find a difficulty imbalance in rejection-sampled datasets and show that more extensive training on difficult samples improves model performance.
@@ 3 Problem setup
Given a vocabulary set $W$, we consider a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{M}$, where $x^{(i)}\in W^{n_{i}}$ is a prompt, $y^{(i)}\in W^{m_{i}}$ is its ground-truth completion, and $n_{i}$ and $m_{i}$ are the prompt and completion lengths. In the context of math, $x^{(i)}$ is the problem statement and $y^{(i)}$ is its solution. To model the conditional distribution $p(y^{(i)}|x^{(i)})$ we use an autoregressive transformer model [39], which is traditionally trained by minimizing the cross-entropy loss
@@
To use and evaluate the model at test time, we assume the existence of an efficient oracle verifier $V$ which takes as input an $(x,y)$ pair and returns $V(x,y)=1$ if $y$ is a correct completion of $x$ and 0 otherwise. Practical examples of verifiers include compilers or pre-defined unit tests for coding problems, or automatic proof checkers in mathematical theorem proving. In such applications, a simple method, known as pass@N, for trading test-time compute for accuracy involves sampling $N$ completions from $\hat{p}$ given the test prompt $x$ and applying the verifier $V$ to all of them to search for a correct solution. The probability of a correct answer is then no longer the probability that $1$ completion is correct, but rather the probability that at least one of $N$ is correct. This probability, for a dataset $\mathcal{D}$, is given by the pass@N coverage metric
$\mathcal{C}_{\mathcal{D}}^{N}=\mathop{\mathbb{E}}_{x\sim\mathcal{D},y_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\hat{p}(\cdot|x)}\mathbb{P}(\exists j\in[N]\,s.\,t.\,V(x,y_{j})=1).$ (2)
@@ 4.1 The CE loss induces overfitting for pass@N
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base [40] on the MATH [41] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training. We follow Lightman et al. [22] and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer without a reasoning trace. We will discuss training with CoT in Section 5.4. We confirm that the results presented in this section also hold for a larger model, Llama-3-70B-base, and direct readers to Section C.3 for additional results.
Table 1: Pass@N coverage on the MATH test set for a Llama-3-8B-base model fine-tuned with CE loss on direct answers from the MATH training set. While pass@1 improves with continued training, for large $N$, surprisingly, pass@N decreases with number of training epochs.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
However, Lemma 4.1 can fail when there is more than one problem. To understand this in a simple setting, consider a test set with two problems $x_{1}$ and $x_{2}$ with unique correct answers $y_{1}$ and $y_{2}$. Consider two models. The first model assigns probabilities $\hat{p}_{1}(y_{1}|x_{1})=1$ and $\hat{p}_{1}(y_{2}|x_{2})=0$. If we think of the probability a model assigns to an answer as its confidence in that answer, this model is highly confident and correct for problem $x_{1}$, but highly confident and wrong for $x_{2}$. Its pass@1 coverage is thus 50%. Moreover, since it always gets $x_{1}$ right and $x_{2}$ wrong, its pass@N coverage remains at 50%. In contrast, consider a second model which assigns probabilities $\hat{p}_{2}(y_{1}|x_{1})=\hat{p}_{2}(y_{2}|x_{2})=0.1$. This model has high confidence (0.9) but unfortunately on incorrect answers. Therefore its pass@1 accuracy is only 10%. However, it is willing to explore or hedge by placing some low confidence ($0.1$) on the other answer, which happens to be correct. Thus if this model samples $N$ times, the pass@N coverage increases with $N$, eventually approaching 100% as $N\rightarrow\infty$. Thus the first model outperforms the second in terms of pass@1 but not pass@N for large $N$. This indicates a tradeoff amongst policies: those that do better on pass@1 may not do better at pass@N. Of course the best one can do is be confident and correct, corresponding to the optimal model with $\hat{p}^{*}(y_{1}|x_{1})=\hat{p}^{*}(y_{2}|x_{2})=1$. However, this toy example reveals that if one cannot guarantee correctness, it can be beneficial to limit confidence and explore more solutions.
@@
This lower bound is a monotonically decreasing function of $N$ (Figure 11), implying that at small $N$, optimal policies must have high max confidence, thereby favoring exploitation and discouraging exploration. Note in the limit $N\to 1^{+}$, the lower bound is always $1$, recovering the intuitive result that the optimal policy for pass@1 is to place 100% confidence on the highest accuracy answer.
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
Figure 1: A model trained with CE loss becomes overconfident in its greedy completions, which harms its pass@N coverage; our proposed DCO objective limits this overconfidence. We fine-tune a Llama-3-8B base model on the MATH dataset to produce direct answers. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when choosing the most likely token at each sampling step. (a) The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. (b) At the end of training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up $N$. (c) Same as (a) but shown for the DCO loss with $N^{\prime}=256$. The model trained on DCO shows a much milder overconfidence effect. (d) The confidence distribution of greedy completions after four epochs with DCO for various choices of $N^{\prime}$. As $N^{\prime}$ increases, the model’s confidence in its greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
@@
Lemma 4.2 suggests that at large $N$ it is beneficial to unconfidently explore by assigning low model confidence to many answers, while Lemma 4.3 suggests that at small $N$ it is beneficial to confidently exploit by assigning high model confidence to one or a few answers. These lemmas make a prediction that could explain the empirical observation in Table 1: namely, maximization of pass@1 makes the model overconfident, thereby preventing good performance on pass@N.
@@
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas 4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure 1 (a). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure 1 (b) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance. In Section C.4, we observe a similar trend of overconfidence on the out-of-distribution test set AIME24, further highlighting the generality of this failure mode.